Donna is 12 years younger than her brother. In 2 years, she will be 14 years more than half as old as her brother.How old is each now?

Answer by checkley79(3341)

D=B-12
D+2=B/2
SUBSTITUTE (B-12) FOR D IN THE SECOND EQUATION.
B-12+2=B/2
B-10=B/2 CROSS MULTIPLY.
B=2(B-10)
B=2B-20
B-2B=-20
-B=-20
B=20 ANS. FOR BROTHER'S AGE.
D=20-12=8 AGE OF DONNA.
PROOF:
8+2=20/2
10=10